twin city the probability that a power engine from CDA (Capital development Authority) is available when needed is 0.45 and the probability that a power engine from RDA (Rawalpindi development Authority) is available when needed is 0.45. What is the probability that the engines neither from CDA nor from RDA are available when needed?
1
Expert's answer
2020-09-28T20:42:13-0400

Solution


let "P(C_a)= 0.45" be the probability that CDA is available. The probability that CDA is unavailable, "P(C_n) =1-0.45=0.55"


Similarly, let "P(R_a)=0.45" be the probability that RDA is available. The probability that RDA is unavailable, "P(R_n) =1-0.45=0.55"


the probability that the engines neither from CDA nor from RDA are available when needed:


"P(C_n)\ and\ P(R_n)""0.55*0.55= 0.3025"

answer: the probability that neither is available is 0.3025